jasmine99:

Donna is taller than David but shorter than Bill. Dan is shorter than Bob. What is the first letter in the name of the tallest person?

Vocaloid:

By definition the tallest person in a group cannot be shorter than anybody else Re-read the problem, and see which people are not described as being shorter than another person in the group

Vocaloid:

Donna is taller than David (so David is shorter than Donna) Donna is shorter than Bill Dan is shorter than Bob So David, Donna, and Dan are all shorter than some other person in the group. So none of these people can be tallest. Who is left?

jasmine99:

Bill or Bob I was thinking

Vocaloid:

Good, either Bill or Bob is the tallest It’s only asking for the first letter of the tallest persons name, and both these peoples names start with B. So it must be B.

jasmine99:

Okay thank you for the help!

jasmine99:

Can I ask one more question I already solved it but just to look over it if anyone can.

Vocaloid:

Sure

jasmine99:

I am thinking of a positive number.  If I square it, double the result, take half of that result, and then add 12, I get 21.  What is my number? For this problem I began by guessing numbers that would work for x and then put them into the problem.  x^2*2/2+12=21 x^2+12=21 x^2=9 x=3

Vocaloid:

Yup that’s correct

jasmine99:

Thank you!